OpenStudy (anonymous):

what is the vertex of the parabola y=x^2-6x+7?

OpenStudy (anonymous):

using completing the square, transform the given equation into vertex form:\[y=a(x-h)^{2}+k\] where (h,k) is the coordinates of the vertex.

OpenStudy (luigi0210):

Using derivatives you get: y'=2x-6 So it's at x=3 :3

OpenStudy (luigi0210):

Data's way works better xD

OpenStudy (anonymous):

Luigi's way is much easier:)

OpenStudy (anonymous):

wait iv seen the h,k thing before but i dont get what you mean?

OpenStudy (anonymous):

check this for further understanding: http://www.mathsisfun.com/algebra/completing-square.html

OpenStudy (campbell_st):

an easy was is the find the line of symmetry for the parabola... as the vertex is on the line of symmetry \[y = ax^2 + bx + c\] the line of symmetry is \[x =\frac{-b}{2a}\] in your question b = -6 and a = 1 substitute them to find the equation and them substitute the value of x you found into the original equation to get the ordered pair for the vertex. is a little easier than factoring to a perfect square

OpenStudy (anonymous):

oh my gosh that was perfect thankyou will do! thanks everyone!